# MSI-NeRF: Linking Omni-Depth with View Synthesis through Multi-Sphere Image aided Generalizable Neural Radiance Field

### v2

@@ 4.1 Implementation Details
For comparison, baseline methods are retrained on the same datasets using the same training strategy. We also keep the network capacity and hyperparameters of them consistent with our implementation.
@@ 4.3 Depth Evaluation
We evaluate the omnidirectional depth estimation performance on a validation set extracted from the OmniHouse and OmniThings. Original and fine-tuned models are used for different datasets, respectively. The qualitative and quantitative evaluation results are shown in Fig. 3 and Tab. 1. We use inverse depth value as our comparison metric, and the unit is set to $m^{-1}$. The qualifier ”$>k$” represents the pixel ratio whose MAE is larger than threshold $k$.
@@ 4.4 Novel View Synthesis Evaluation
Method Generalizable PSNR SSIM LPIPS
MatryODShka 29.665 0.584 0.419
NeRF-360-NGP 28.594 0.451 0.541
Ours 37.302 0.957 0.077
@@ 4.5 Ablation Studies
Although generating good depth maps, the method without additional color input struggles at color learning, producing blurry images. Method without radiance field rendering directly outputs color and occupancy spheres, causing ghosting effects in the synthesized views. Method without separated decoders performs well at novel view synthesis tasks but lacks the ability to estimate fine-grained depth maps. These results show that both designs play an essential role in our method. Ours full model uses color projection input to gain additional color information and implicitly mix them to achieve more reasonable color output. The usage of radiance field rendering creates a more consistent image generation and results in finer color and depth generation. Moreover, the design of separated geometry and appearance decoder makes the role of the network clearer and avoids data confusion. These ablation experiments prove the rationality and innovation of our framework.
@@ 4.6 Real-world Experiment
To test our approach’s generalizability and real-world performance, we constructed a self-collected dataset using a multi-fisheye camera rig with similar settings to the synthetic datasets. The data capture system, fisheye image samples, and the generated results can be seen in Fig. 8. From the generated depth and novel view images, we can see that our method can be well generalized from sim to real. The real-world experiment also demonstrates the practical application value of this method for virtual reality video shooting and 3D scene reconstruction in real-world scenes. However, our method has limitations because it requires accurate camera intrinsic and extrinsic parameters and calibration errors may cause blurring and artifacts in color and depth images in real-world scenes.
5 Conclusion
@@
This paper explores the problem of generating a panoramic scene representation using only four fisheye images as input. To solve this, our presented MSI-NeRF combines the advantage of multi-sphere image representation with volume rendering to construct a generalizable NeRF. From such representation, accurate omnidirectional depth estimation and high-quality novel view synthesis can be conducted efficiently, with only a single network inference. As a result, our method outperforms baseline methods in both image quality and feasibility, which is of great significance for virtual reality and robotics applications.
6 Additional Implementation Details
Since MatryODShka [1] has different input and output format to our method, to achieve fair comparison, we re-implement MatryODShka and match its data format with ours. MatryODShka uses two omnidirectional stereo (ODS) images as input and sphere sweep the images as a sphere sweep volume (SSV). We find that four fisheye images can provide larger FoV than the original two ODS images, so we directly use fisheye images to construct the SSV. For the network MatryODShka used to process SSV, we use the same 3D CNN structure as our MSI decoder by just changing the input and output channel number. We modify the output of the network to generate a blending weight using soft max layer. Alpha blending is then conducted and the final color and depth images can be rendered.
7 Network Architecture
We basically follow the CNN structure proposed in [38] and modify the output format to connect it with our NeRF [29] MLP. The CNN network channel is declined to save GPU memory for the additional MLP and neural rendering. The input of the network is four fisheye images. We use the residual convolution blocks for the image feature extraction, and the dilated convolution for the larger receptive field. The output feature map size is half of the input image. We use a shallow MLP that takes coordinate, direction, projected color and interpolated feature vector as input. The output color and occupancy are then rendered into color and depth map.
8 Derivation Details
We give a detailed derivation of Equation (6) in the main body. Given the ray origin $\mathbf{r}_{o}=(o_{x},o_{y},o_{z})^{T}$ and ray direction $\mathbf{r}_{d}=(d_{x},d_{y},d_{z})^{T}$, the equation of the ray can be given as:
$\mathbf{r}(z)=\mathbf{r}_{o}+z\mathbf{r}_{d},$ (12)
where z is the parameter indicating the distance along the ray, the depth. Since the sphere equation can be writen as $x^{2}+y^{2}+z^{2}=d^{2}$, we substitute $\mathbf{r}(z)$ into the sphere equation:
$\left(r_{o_{x}}+zr_{d_{x}}\right)^{2}+\left(r_{o_{y}}+zr_{d_{y}}\right)^{2}+\left(r_{o_{z}}+zr_{d_{z}}\right)^{2}=d^{2}.$ (13)
We can simplify it to:
$r_{o_{x}}^{2}+r_{o_{y}}^{2}+r_{o_{z}}^{2}+2z\left(r_{o_{x}}r_{d_{x}}+r_{o_{y}}r_{d_{y}}+r_{o_{z}}r_{d_{z}}\right)+z^{2}=d^{2}.$ (14)
From this, we can tell that the problem is to solve a quadratic equation and the coefficients of the quadratic term, linear term, constant term can be represented as:
$a=\mathbf{r}_{d}^{T}\mathbf{r}_{d},\quad b=2\mathbf{r}_{d}^{T}\mathbf{r}_{o},\quad c_{n}=\mathbf{r}_{o}^{T}\mathbf{r}_{o}-1/\left(d^{-1}\right)^{2}.$ (15)
Then the original formula can be established.
9 Application Prospects
Being able to synthesize novel views in 6DoF and estimate depth map panoramically mean a lot to VR and robotics.
Traditional VR contents are usually shot by image stitching, which is simply a stream of panorama images. This results in a limited 3DoF camera movement which can cause omitted parallax information, low immersion, and VR sickness. Instead, our method retains the parallax information within the multi-camera system and achieves 6DoF rendering. This can bring a whole new experience to VR content without changing the shooting equipment. Further more, different from the existing panoramic cameras that only support three-degree-of-freedom re-movement and editing, the video captured by this method can be moved and edited in the translation direction. This is of great significance for filming and video making.
On the robotics side, being able to produce omnidirectional depth estimation also helps a lot. It bring perception in autonomous robots to a new level. Compared to LiDAR, it is much cheaper, with wider FoV, and denser points per frame. The novel view synthesis ability also makes driving cars or drones under a third person’s view possible, which benefits a lot in teleoperation and navigation.
10 Limitations
Limited by the memory size of the GPU device, our network capacity is restricted and the resolution of the generated color and depth map is set to $512\times 256$. Also, due to the limited computation performance, our method can only run at around three frames per second. These limitations makes our method hard to applied in applications with strong real-time requirements.
At the same time, our method also requires accurate extrinsic and intrinsic calibration. These parameters are used in the warping process that turns fisheye projection into MSI representation, which is essential to the final result. In our real-world experiment, the generated color and depth image both suffer from poorly calibrated cameras, which results in blurry and artifacts. The quality of real-world generation will degrade compared to the results from synthetic data.
Due to the usage of MSI, the free roaming of the camera is limited inside the smallest sphere. This way, the setting of the smallest radius can be a trade of between the nearest collision limit and the virtual camera moving range. A bigger inner sphere can lead to a wider moving range, while may cause objects close to camera being cut off. This trade off needs careful tuning when faced with different scenarios.
11 Additional Experiments Result
Here, we show the additional generated depth estimation and novel view synthesis results on the OmniHouse dataset. They are shown in Fig. 9. More visualization results can be seen in the video in the supplementary files.
Figure 9: Additional experiment results.

### v3

@@ 4.3 Depth Evaluation
We evaluate the omnidirectional depth estimation performance on a validation set extracted from the OmniHouse and OmniThings. Original and fine-tuned models are used for different datasets, respectively. We reimplement a compressed version of OmniMVS [38] using the same backbone as ours, keeping the network capacity and hyperparameters consistent with ours. The qualitative and quantitative evaluation results are shown in Fig. 3 and Tab. 1. We use inverse depth value as our comparison metric, and the unit is set to $m^{-1}$. The qualifier ”$>k$” represents the pixel ratio whose MAE is larger than threshold $k$.
@@ 4.4 Novel View Synthesis Evaluation
Method Generalizable PSNR SSIM LPIPS Time
MatryODShka 29.665 0.584 0.419 0.128s
NeRF-360-NGP 28.594 0.451 0.541 $\sim$10s
Ours 37.302 0.957 0.077 0.383s
@@ 4.5 Ablation Studies
The method without additional color input struggles at color learning, producing blurry images. The method without radiance field rendering directly outputs color and occupancy spheres, causing ghosting effects in the synthesized views. The method without separated decoders lacks the ability to estimate fine-grained depth maps. Ours full model uses color projection input to gain additional color information and implicitly mix them to achieve more reasonable color output. The usage of radiance field rendering creates a more consistent image generation and results in finer color and depth generation. Moreover, the design of separated geometry and appearance decoder makes the role of the network clearer and avoids data confusion.
@@
Figure 8: Visualization of our fisheye image input with regular image input.
To highlight the advantage of our fisheye input format, we compare our method against the method that uses regular image input. We reimplement our method using regular image input by warping the input to the MSI structure. The result shows that more than 16 regular images are needed to achieve similar reconstruction quality.
@@ 4.6 Real-world Experiment
To test our approach’s generalizability and real-world performance, we constructed a self-collected dataset using a multi-fisheye camera rig with similar settings to the synthetic datasets. The data capture system, fisheye image samples, and the generated results can be seen in Fig. 9. From the generated depth and novel view images, we can see that our method can generalize from sim to real. The real-world experiment also demonstrates the practical application value for virtual reality video shooting and 3D reconstruction in real-world scenes. However, limitations exist because it requires accurate camera intrinsic and extrinsic parameters and calibration errors may cause blurring and artifacts in color and depth images in real-world scenes.
5 Conclusion and Limitation
@@
This paper explores the problem of generating a panoramic scene representation using only four fisheye images as input. Our presented MSI-NeRF combines the advantage of MSI representation with volume rendering to construct a generalizable NeRF. From such representation, accurate omni-depth estimation and high-quality novel view synthesis can be conducted efficiently, with only a single network inference. Our method outperforms baseline methods in both image quality and feasibility, which is significant for virtual reality and robotics applications. Since our method uses the idea of MVS and does not have generation ability, the reconstruction of the occluded part is unpleasant. In the future research, we hope to leverage the power of generative models to inpaint the unseen part.